# Generative Topological Networks

### v2

@@ Abstract
Generative models have seen significant advancements in recent years, yet often remain challenging and costly to train and use. We introduce Generative Topological Networks (GTNs) – a new class of generative models that addresses these shortcomings. GTNs are trained deterministically using a simple supervised learning approach grounded in topology theory. GTNs are fast to train, and require only a single forward pass in a standard feedforward neural network to generate samples. We demonstrate the strengths of GTNs on several datasets, including MNIST, CelebA and the Hands and Palm Images dataset. Finally, the theory behind GTNs offers insights into how to train generative models for improved performance. Code and weights are available at: https://github.com/alonalj/GTN
@@ 1 Introduction
Figure 1: Samples generated by a GTN trained on CelebA $64\times 64$ with a latent dimension of $100$.
@@
Deep generative models such as Generative Adversarial Networks (GANs) (Goodfellow et al. 2020), Variational Autoencoders (VAEs) (Kingma and Welling 2013), Energy-Based Models (EBMs) (LeCun et al. 2006; Ngiam et al. 2011), normalizing flows (NFs) (Rezende and Mohamed 2015), and diffusion models (Sohl-Dickstein et al. 2015b), have demonstrated remarkable capabilities for generating samples based on training data distributions (Kang et al. 2023; Ho, Jain, and Abbeel 2020; Ramesh et al. 2021; Ho et al. 2019; Kingma and Dhariwal 2018; Kingma et al. 2016; Reed et al. 2017; Van Den Oord, Vinyals et al. 2017; Ramesh et al. 2021; Onken et al. 2021).
Diffusion models, in particular, have seen great advancements over the last few years in a variety of settings (Ho, Jain, and Abbeel 2020; Luo et al. 2023; Nichol and Dhariwal 2021; Dhariwal and Nichol 2021; Ho et al. 2022; Rombach et al. 2022), and possess several advantages compared to other methods. For example, compared to GANs and VAEs diffusion models have more training stability as they do not suffer from mode collapse like GANs (Zhang, Li, and Yu 2018), or posterior collapse like VAEs (Lucas et al. 2019).
However, diffusion models have several downsides. Most prominent is their inferior time-efficiency during training and sampling when compared to VAEs and GANs. This is because they employ a gradual and stochastic transition from a Gaussian distribution to the data distribution, performed over multiple time-steps known as the reverse diffusion process. The complexity of this process also makes it challenging to produce interpolations – desirable in the context of continuous data like audio and video. Although recent advancements have made significant steps towards enhancing diffusion models, the complex diffusion process is still a key challenging component (Manduchi et al. 2024).
@@
In this work we introduce a particularly efficient, stable and simple class of generative models – Generative Topological Networks (GTNs). GTNs learn a continuous and invertible mapping between two distributions in a supervised manner that facilitates efficient training and sampling. Specifically, given a training set of samples (e.g. images) and a tractable source distribution (e.g. Gaussian), GTNs learn a function $\hat{h}$ such that, given a $y$ sampled from the source distribution, $\hat{h}(y)$ is a sample representing the training data distribution. GTNs are reminiscent of NFs, which aim to transform one distribution into another using a sequence of invertible and differentiable maps. GTNs, however, are constructed entirely differently, and offer many advantages from both practical and theoretical perspectives.
@@
From a practical perspective, GTNs are extremely simple and efficient, requiring a single, vanilla, feedforward architecture trained using standard supervised learning. This means that GTNs avoid the complexities of introducing stochasticity during training which is present in VAE and diffusion; circumvent the intricacies of training more complex architectures such as those employed by diffusion and GANs; are not a likelihood-based approach, as are most generative models; and do not pose constraints on the structure of the neural network, as do NFs. GTNs also operate on latent representations of the data, providing even more enhanced training and sampling efficiency. These advantages manifest in our experiments – realistic samples are obtained at early training epochs, and generating a single image takes around $4$ milliseconds on a standard PC and $0.003$ milliseconds on a T4 GPU.
@@
From a theoretical perspective, GTNs provide guarantees and properties that are desirable in the context of generative models. These include: learnability (via the universal approximation theorem), continuity (for continuous interpolations – Figure 6), bijectivity (for diversity and coverage of the data distribution – Figure 3) and properties that serve as guiding topological principles on how to train generative models for better performance (see the Method section and the swiss-roll example).
The remainder of the paper is structured as follows: In the Method section we first develop the theory behind GTNs for the $1$-dimensional (1D) case and use it to generate samples from the swiss-roll. We then extend this theory to higher dimensions and use it to generate samples from the multivariate uniform distribution. In the Experiments section we apply GTNs to real datasets, also including demonstrations of continuous interpolation. We follow with Related Work section and conclude with the Discussion section.
@@ 2 Method
The Significance of $h$ Being a Homeomorphism.
@@
In this example, we will see that samples from the swiss-roll can be easily generated from a $1$D normal distribution with visibly near-perfect accuracy (Figure 3 (A)) using a GTN. This would not be as easy using a $2$D normal distribution (the approach taken by previous methods such as (Sohl-Dickstein et al. 2015b)), as we will now explain.
The swiss-roll is a $1$D manifold in a $2$D space $\mathbb{R}^{2}$ (generated using a single, 1-dimensional, continuous random variable $\theta$). Being a $1$D manifold, it cannot be homermorphic to the support of a $2$D normal distribution ($\mathbb{R}^{2}$) since homeomorphic manifolds must have the same dimension. Therefore, there would be no hope of learning a homeomorphism that maps a $2$D Gaussian random variable to the swiss-roll. However, the swiss roll is homeomorphic to $\mathbb{R}$ – the support of a $1$D standard normal distribution. In fact, since $\theta$ is a random variable that satisfies the conditions in Theorem 2.1, it is homeomorphic to $\mathbb{R}$ via $h$, which we can learn using $\hat{h}$. The fact that there is no hope of learning a homeomorphism in a higher-dimensional space than the intrinsic dimension is also important for NFs, which use a diffeomorphism (a type of homeomorphism) as a generator, and it may also help explain why latent diffusion models have shown improvements over pixel-space diffusion models (Rombach et al. 2022).
To train $\hat{h}$, we created a dataset of $n=50,000$ samples from $\theta$, denoted $D_{X}$. We sampled $n$ samples from a $1$D standard normal distribution to obtain $D_{Y}$. We labeled each $y_{i}\in D_{Y}$ with its $x_{y_{i}}\in D_{X}$ as defined in the previous section. We used a standard feedforward neural network as $\hat{h}$ ($4$ layers of width $6$; details in Appendix Table 1). Figure 3 (A) shows the result of testing the trained model $\hat{h}$ on a set of new samples $y_{1},\dots,y_{k}$ drawn from $\mathcal{N}(0,1)$ (each point is obtained by predicting $\hat{\theta}_{i}:=\hat{h}(y_{i})$ and applying the formula for the swiss roll to $\hat{\theta}_{i}$).
For comparison, we refer the reader to Figure 1 (middle) in (Sohl-Dickstein et al. 2015b) where a trained diffusion model was applied to points sampled from a standard $2$D normal distribution. The points are gradually and stochastically moved back towards the swiss-roll manifold, but the result contains points that fall outside of this manifold. With $\hat{h}$, the generated points lie entirely within the data manifold. In fact, this holds true early in the training process, as shown in Appendix Figure 8.
@@ 2.2 Higher Dimensions
For more complicated distributions, it may be difficult to explicitly define a similar $h$ since $h_{1}$ would depend on the line. Nevertheless, the simpler case above can guide us on how to train a neural network to learn one (as accomplished in Figure 3 (B)).
@@
Figure 5: Randomly chosen samples generated by GTN during the training process on MNIST at four points of improvement on the validation set. Each epoch shows 100 images, each of which is the decoded $\hat{h}(r)$ for some random $r\sim\mathcal{N}(\textbf{0},\textbf{I})$). The rightmost grid contains 100 images each of which is the decoded latent vector of a random real image.
@@ 3 Experiments
In light of the discussion in Section 2.1 on using the intrinsic dimension (ID) of the data, we use autoencoders to represent each dataset in a lower dimensional latent space before training $\hat{h}$. Each dataset was trained separately (with its own autoencoder and $\hat{h}$). In all experiments, we use the same vanilla autoencoder architecture adapted to different latent dimensions (the encoder consisted of $2$ convolution layers with ReLU activation followed by a fully connected layer). For $\hat{h}$ we use a standard feedforward neural network with the width and depth depending on the data. All architecture and training details for all experiments are in Appendix F.
@@
To train $\hat{h}$, we set $X$ to be the latent vectors of the training set (the encoded images), and set $Y$ to be the standard multivariate normal distribution of the same dimension as the latent space (e.g., if we used a latent dimension of $d=100$ in the autoencoder, then $Y$ has dimension $100$). Images were generated using Algorithm 2 by computing: $autoencoder.decoder\big{(}\hat{h}(r)\big{)}$ where $r\sim\mathcal{N}(\textbf{0},\textbf{I})$ (generation occurs in the latent space).
Results.
We first apply our method to MNIST (LeCun et al. 1998) using a latent dimension of $5$. We chose $d=5$ to allow for comparison with VAE for the same latent dimension (Figure 5(b) in (Kingma and Welling 2013)). Figure 5 shows random sets of generated images for MNIST at several epochs of improvement on the validation set, as well as a sample of random real images for comparison. These results show that the generated samples were already realistic even after one epoch (a similar observation was made for the swiss-roll). Besides the aforementioned comparison to VAE, further comparisons can be made with diffusion (Figure App.1. in (Sohl-Dickstein et al. 2015b)) and a continuous normal flow (CNF) (Figure 3(a) (right) in (Huang and Yeh 2021)). Training was stopped when no improvement in the validation MSE was obtained for $3$ epochs.
@@ Results.
Next we applied the method to CelebA (Liu et al. 2015) (Figure 1 and Appendix Figures 10-14). Images were center cropped to $148\times 148$ and resized to $64\times 64$ We tested latent dimensions of $d\in\{10,25,50,100\}$, which are within the range of the typical ID estimated for image datasets ($10$-$50$) or close ($100$) (Pope et al. 2021). We used Inception Score (IS) (Salimans et al. 2016), a common evaluation metric, as the stopping criteria: after every epoch, the GTN generated $200$ random images for which the IS was evaluated. We used IS instead of the validation set since the IS continued to improve well beyond the point of plateau on the validation set (for MNIST the IS was irrelevant since the data is grayscale). Training stopped after $300$ epochs of no improvement in IS. We also produced the IS for a single random set of $200$ real decoded images as a baseline (”real IS”). Plotting the IS by epoch, alongside the real IS (Appendix Figure 9) shows that the real IS is either achieved ($d=10,25$) or nearly achieved ($d=50,100$) by the GTN and that the gap between IS and real IS decreases as $d$ decreases. With the final model, we also computed another common evaluation metric – the Fréchet Inception Distance (FID) (Heusel et al. 2017) – between the training set and an equally-sized set of generated images. We observed a similar pattern, namely that FID decreased as $d$ decreased, except for $d=10$) (Appendix Table 2). Overall, the IS and FID, as well as the difference between IS and real IS, improved as $d$ decreased. Visual inspection of randomly generated images next to real images across all dimensions (Appendix Figures 10-13) confirms these improvements. Namely, as $d$ decreases we see more similarity between generated images and real images as well as improved diversity.
@@
Observing the progress in image generation during training (Appendix Figure 14), we can see that realistic images were already obtained at half the training time (see epoch $276$). One epoch took just under $1$ minute ($50.6$ seconds on average) on a single T4 GPU, reflecting $9$ hours until the last improvement in IS at epoch $640$ and $<4$ hours to reach epoch $276$. The time to generate $1,000$ samples (i.e. to compute: $autoencoder.decoder\big{(}\hat{h}(r)\big{)}$) on a standard PC was $4.07$ seconds (reflecting $4.07$ milliseconds per image) and $0.00367$ seconds (reflecting $0.00367$ milliseconds per image) on a T4 GPU.
@@
Figure 6 shows interpolations for both CelebA and HaP. For CelebA we used the $d=100$ model and for HaP we used $d=50$. The dimensions were chosen after consulting the IS and FID results across all four $d$ values in Appendix Table 2 while preferring a higher dimension ($d\in\{50,100\}$) to retain more visual detail (e.g., compare Appendix Figures 10-11 to 12-13): in CelebA $d=100$ obtained similar results to $d=50$ and in HaP $d=50$ obtained better results than $d=100$. Note that we observe similar trends in HaP as in CelebA: the difference between IS and real IS, as well as the FID decrease with $d$ (Appendix Figure 16 and Table 2).
Figure 6 shows that GTN indeed generates continuous interpolations for both CelebA and HaP. Each row contains generated interpolations between two images (leftmost and rightmost). Specifically, each row contains the results of decoding $\hat{h}\big{(}\lambda y_{\text{left}}+(1-\lambda)y_{\text{right}}\big{)}$ for $20$ linearly spaced $\lambda\in[0,1]$, where, in CelebA, both $y_{\text{left}}$ and $y_{\text{right}}$ were sampled from $\mathcal{N}(\textbf{0},\textbf{I})$, and in HaP $y_{\text{left}}$ and $y_{\text{right}}$ are each the normal label of a real encoded image (from Algorithm 1), chosen so that they have the same orientation.

### v3

@@ Abstract
Generative methods have recently seen significant improvements by generating in a lower-dimensional latent representation of the data. However, many of the generative methods applied in the latent space remain complex and difficult to train. Further, it is not entirely clear why transitioning to a lower-dimensional latent space can improve generative quality. In this work, we introduce a new and simple generative method grounded in topology theory – Generative Topological Networks (GTNs) – which also provides insights into why lower-dimensional latent-space representations might be better-suited for data generation. GTNs are simple to train – they employ a standard supervised learning approach and do not suffer from common generative pitfalls such as mode collapse, posterior collapse or the need to pose constraints on the neural network architecture. We demonstrate the use of GTNs on several datasets, including MNIST, CelebA, CIFAR-10 and the Hands and Palm Images dataset by training GTNs on a lower-dimensional latent representation of the data. We show that GTNs can improve upon VAEs and that they are quick to converge, generating realistic samples in early epochs. Further, we use the topological considerations behind the development of GTNs to offer insights into why generative models may benefit from operating on a lower-dimensional latent space, highlighting the important link between the intrinsic dimension of the data and the dimension in which the data is generated. Particularly, we demonstrate that generating in high dimensional ambient spaces may be a contributing factor to out-of-distribution samples generated by diffusion models. We also highlight other topological properties that are important to consider when using and designing generative models. Our code is available at: https://github.com/alonalj/GTN
@@ 1 Introduction
Figure 1: Samples generated by a GTN trained on a latent representation of CelebA $64\times 64$ with latent dimension $d=100$.
@@
Deep generative models such as Generative Adversarial Networks (GANs) (Goodfellow et al. 2020), Variational Autoencoders (VAEs) (Kingma and Welling 2013), Energy-Based Models (EBMs) (LeCun et al. 2006; Ngiam et al. 2011), normalizing flows (NFs) (Rezende and Mohamed 2015), and diffusion models (Sohl-Dickstein et al. 2015a), have demonstrated remarkable capabilities for generating samples based on training data distributions (Kang et al. 2023; Ho, Jain, and Abbeel 2020; Ramesh et al. 2021; Ho et al. 2019; Kingma and Dhariwal 2018; Kingma et al. 2016; Reed et al. 2017; Van Den Oord, Vinyals et al. 2017; Ramesh et al. 2021; Onken et al. 2021).
Many image-generation methods like diffusion are often applied in pixel-space (Ho, Jain, and Abbeel 2020; Sohl-Dickstein et al. 2015b). More recently, generative methods have seen great improvements in generative quality by transitioning to a lower-dimensional latent representation of the data (Rombach et al. 2022). Although training in a lower-dimensional latent space reduces the computational burden, the methods used to generate in the latent space remain complex and computationally expensive. Further, besides the obvious training efficiency, it is not entirely clear why transitioning to a lower-dimensional latent representation of the data leads to improved generative quality compared to pixel-space, or by how much the dimension should be reduced.
In this work we introduce a new class of generative models – Generative Topological Networks (GTNs) – that provides a simple approach for generating in the latent space and that sheds light on why transitioning to a lower-dimensional latent space improves generative quality.
@@
We begin by introducing GTNs. Given a training set of samples (e.g. images) and a tractable source distribution (e.g. Gaussian), GTNs learn to approximate a continuous and invertible function $h$ such that, given a $y$ sampled from the source distribution, $h(y)$ is a sample representing the training data distribution. GTNs are reminiscent of NFs, which aim to transform one distribution into another using a sequence of invertible and differentiable maps. NFs, however, pose specific constraints on the network architecture. In contrast, GTNs do not pose any constraints, and are fully operational using a simple vanilla feedforward architecture.
@@
From a practical perspective, GTNs are simple to train, requiring only a single, vanilla, feedforward architecture trained using standard supervised learning. This allows GTNs to avoid issues like mode collapse or posterior collapse faced by GANs and VAEs; to circumvent the intricacies of training more complex architectures such as those employed by diffusion and GANs; and to avoid posing constraints on the structure of the neural network, as in NFs. These advantages manifest in our experiments – with realistic samples obtained at early epochs using a single T4 GPU.
@@
From a theoretical perspective, GTNs provide guarantees and properties that are desirable in the context of generative models. These include: learnability (via the universal approximation theorem), continuity (for continuous interpolations – Figure 8), bijectivity (for diversity and coverage of the data distribution – Figure 3) and properties that serve as guiding topological principles that can inform the design of generative methods (see the Method section and the swiss-roll example). As an example, we use some of the topological principles behind GTN to demonstrate that generating with diffusion without transitioning to a lower-dimensional latent space can lead to many out-of-distribution samples. We also discuss why this is important for other methods too.
The remainder of the paper is structured as follows: In the Method section we first develop the theory behind GTNs for the $1$-dimensional (1D) case and use it to accurately generate samples from an intrinsically 1D swiss-roll represented in a 2D ambient space. Here, we emphasize the significance of the intrinsic dimension of the data for generative models and demonstrate that diffusion struggles to accurately generate the 1D swiss-roll in the original 2D ambient space. We then extend the theory behind GTNs from the 1D case to higher dimensions and use it to accurately generate samples from the multivariate uniform distribution. Finally, we apply GTNs to real datasets represented in a lower-dimensional latent space obtained by autoencoders. We show that GTNs generate samples resembling reconstructed data and that they improve upon VAEs both quantitatively and qualitatively. We conclude with Related Work and Discussion sections where we elaborate on how GTNs compare to other methods from a practical perspective, and demonstrate extensions of GTNs to more complicated data.
@@ 2 Method
On the many advantages of $h$ being a homeomorphism in the context of generative models see Appendix H.
@@
In this example, we show that samples from the swiss-roll can be easily generated from a $1$D normal distribution with visibly near-perfect accuracy (Figure 3 (A)) using a GTN. This would not be as easy using a $2$D normal distribution, as explained next, and as shown by an attempt to do so with diffusion in Figure 4.
Figure 4: Samples generated by a diffusion model when trained using a 2D Gaussian and a 2D representation of the 1D swiss-roll. Many of the generated samples fall out-of-distribution. We use the diffusion model provided by (Jiménez 2024), adapting only the input. The plot format is also by (Jiménez 2024).
The swiss-roll is a $1$D manifold in a $2$D space $\mathbb{R}^{2}$ (generated using a single, 1-dimensional, continuous random variable $\theta$). Being a $1$D manifold, it cannot be homermorphic to the support of a $2$D normal distribution ($\mathbb{R}^{2}$) since homeomorphic manifolds must have the same dimension. Therefore, there would be no hope of learning a homeomorphism that maps a $2$D Gaussian random variable to the swiss-roll. However, the swiss roll is homeomorphic to $\mathbb{R}$ – the support of a $1$D standard normal distribution. In fact, since $\theta$ is a random variable that satisfies the conditions in Theorem 2.1, it is homeomorphic to $\mathbb{R}$ via $h$, which we can learn using $\hat{h}$. The fact that there is no hope of learning a homeomorphism in a higher-dimensional space than the intrinsic dimension is also important for other generative methods like NFs, which use a diffeomorphism (a type of homeomorphism) as a generator, and it may also help explain why latent diffusion models have shown improvements over pixel-space diffusion models (Rombach et al. 2022).
To train $\hat{h}$, we created a dataset of $n=50,000$ samples from $\theta$ (Appendix G), denoted $D_{X}$. We sampled $n$ samples from a $1$D standard normal distribution to obtain $D_{Y}$. We labeled each $y_{i}\in D_{Y}$ with its $x_{y_{i}}\in D_{X}$ as defined in the previous section. We used a standard feedforward neural network as $\hat{h}$ ($4$ layers of width $6$; details in Appendix Table 2). Figure 3 (A) shows the result of testing the trained model $\hat{h}$ on a set of new samples $y_{1},\dots,y_{k}$ drawn from $\mathcal{N}(0,1)$ (each point is obtained by predicting $\hat{\theta}_{i}:=\hat{h}(y_{i})$ and applying the formula for the swiss roll to $\hat{\theta}_{i}$).
For comparison, we refer the reader to Figure 1 (middle) in (Sohl-Dickstein et al. 2015a) where a trained diffusion model was applied to points sampled from a standard $2$D normal distribution. The points are gradually and stochastically moved back towards the swiss-roll manifold, but the result contains points that fall outside of this manifold. With $\hat{h}$, the generated points lie entirely within the data manifold. In fact, this holds true early in the training process, as shown in Appendix Figure 12.
@@ 2.2 Higher Dimensions
For more complicated distributions, it may be difficult to explicitly define a similar $h$ since $h_{1}$ would depend on the line. Nevertheless, the simpler case above can guide us on how to train a neural network to learn one (as accomplished in the non-rotation-invariant 2D uniform case in Figure 3 (B)).
@@
Figure 6: Random generated and reconstructed images for CIFAR-10 (left), CelebA (centre) and MNIST (right). Each of the generated samples is the decoded $\hat{h}(r)$ for a random $r\sim\mathcal{N}(\textbf{0},\textbf{I})$) and each of the reconstructions is the decoded vector of a random real image.
@@ 3 Experiments
In light of the discussion in Section 2.1 on using the intrinsic dimension (ID) of the data, we use autoencoders to represent each dataset in a lower dimensional latent space before training the GTN $\hat{h}$. Each dataset was trained separately (with its own autoencoder and $\hat{h}$). Unless mentioned otherwise, all experiments used the same vanilla autoencoder architecture adapted to different latent dimensions (the encoder consisted of two convolution layers with ReLU activation followed by a fully connected layer). For $\hat{h}$ we use a standard feedforward neural network with the width and depth depending on the data. Architecture and training details can be found in Appendix F and Table 2 with further details in our code.
@@
To train $\hat{h}$, we set $X$ to be the latent vectors of the training set (the encoded images), and set $Y$ to be the standard multivariate normal distribution of the same dimension as $X$ (e.g., for a latent dimension of $d=100$ in the autoencoder, $Y$ has dimension $100$). Images were generated using Algorithm 2 by computing: $autoencoder.decoder\big{(}\hat{h}(r)\big{)}$ where $r\sim\mathcal{N}(\textbf{0},\textbf{I})$ (generation occurs in the latent space).
Evaluation.
We applied our method to MNIST, CIFAR-10, CelebA $64\times 64$ and the Hands and Palm (HaP) datasets.
We first applied our method to MNIST (LeCun et al. 1998) using a latent dimension of $d=5$ to allow for comparison with VAE for the same latent dimension (see Figure 5(b) in (Kingma and Welling 2013)). Figure 6 (right) shows random sets of generated images for MNIST, as well as a sample of random reconstructed images for comparison.
@@
Next we applied the method to CelebA (Liu et al. 2015) (Figure 1 and Appendix Figure 15). Images were center cropped to $148\times 148$ and resized to $64\times 64$. We tested latent dimensions of $d\in\{10,25,50,100\}$, which are within the range of the typical ID estimated for image datasets ($10$-$50$) or close ($100$) (Pope et al. 2021). We used Inception Score (IS) (Salimans et al. 2016), a common evaluation metric, as the stopping criteria: after every epoch, the GTN generated $200$ random images for which the IS was evaluated. We used IS instead of validation-set results since the IS continued to improve well beyond the point of plateau on the validation set. Training stopped after $300$ epochs of no improvement in IS. For evaluation, we used both IS and Fréchet Inception Distance (FID) (Heusel et al. 2017). For IS, in addition to the best IS obtained by the model, we produced the IS for a single random set of $200$ reconstructed images (”recon-IS”) since this reflects the best possible IS that we can reasonably hope to achieve. Plotting IS and recon-IS by epoch, (Appendix Figure 13) shows that IS increases throughout the training process, and that recon-IS is either achieved ($d\in\{10,25\}$) or nearly achieved ($d\in\{50,100\}$) by the GTN. For FID, the lowest FID was for $d=100$ with 66.05 and the highest was for $d=10$ with 119.46. FID and IS across all dimensions are reported in Appendix Table 3.
@@
Observing the progress in image generation during training (Appendix Figure 15), shows that realistic images were already obtained at half the training time (see epoch $276$). One epoch took just under $1$ minute ($50.6$ seconds on average) on a single T4 GPU, reflecting $9$ hours until the last improvement in IS at epoch $640$ and $<4$ hours to reach epoch $276$. The relatively fast convergence raises the question whether the latent space was approximately normal to begin with. Appendix Figure 17, which compares images generated by decoding random normal vectors with images generated using the GTN, demonstrates that this is not the case.
Next, we designed a controlled experiment to compare our method to the closely-related VAE on both CelebA and CIFAR-10. We used the vanilla VAE suggested for CelebA from (AntixK 2024) adapting only the latent dimension, and trained two autoencoders: (1) The suggested VAE and; (2) A vanilla autoencoder later used to train a GTN on the learned latent representations. The transition from VAE to vanilla autoencoder included only the necessary adaptations, so that all remaining architectural considerations and hyper-parameters were identical between the two autoencoders. We then used the latent representations from the vanilla autoencoder to train a GTN, where the GTN had the same architecture as in the earlier experiments. No tuning of any kind was performed, including of the random seed (which was set once at the very start of the entire project). Particularly, the GTN architecture and training settings (for both CelebA and CIFAR-10) were unchanged from the previous CelebA experiments (except for the dimensions of the input, output and latent space). Figure 7 shows a side-by-side comparison of randomly generated images from both methods. Table 1 provides the FID and IS results, including for reconstruction quality (i.e. recon-FID and recon-IS, which compare reconstructions to real images). A comparison with reconstructions is also shown in Figure 6. Although GTN outperforms VAE on both CelebA and CIFAR-10, the CIFAR-10 results are overall worse compared to the CelebA results. This is likely since: (a) the autoencoder settings were originally suggested for CelebA; and (b) no fine-tuning or architecture search was performed for the GTN when transitioning from the CelebA experiment to CIFAR-10. For additional context with other methods, we also provide Appendix Tables 4 and 5.
Figure 7: Controlled comparison of GTN with VAE (Kingma et al. 2016) on CelebA and CIFAR-10 – randomly generated images from both.
Data Method IS $\uparrow$ recon-IS $\uparrow$ FID $\downarrow$ recon-FID $\downarrow$
CelebA VAE 1.0 1.0 94.66 83.48
GTN 1.90 1.94 73.18 67.08
CIFAR-10 VAE 1.77 1.90 286.49 277.48
GTN 2.07 2.24 238.62 181.53
Table 1: Controlled comparison of GTN with VAE (Kingma et al. 2016) on CelebA and CIFAR-10. The VAE result was obtained by training the vanilla VAE architecture supplied in (AntixK 2024). Prior to training the GTN we trained a new autoencoder that employs the same architectures as the vanilla VAE, with only the necessary adaptations to transition from VAE to a vanilla autoencoder. GTN is trained on the latent representations provided by this autoencoder, with $d=100$ and $d=128$ for CelebA and CIFAR-10, respectively. We use 10,000 random training images and an equal number of random generated samples (decoded generated latent vectors) to compute FID. We use 200 random generated and 200 random real images for IS.
@@
Figure 8 shows that GTN indeed generates continuous interpolations. Each row contains generated interpolations between two images (leftmost and rightmost). Specifically, each row contains the results of decoding $\hat{h}\big{(}\lambda y_{\text{left}}+(1-\lambda)y_{\text{right}}\big{)}$ for $20$ linearly spaced $\lambda\in[0,1]$, where, in CelebA, both $y_{\text{left}}$ and $y_{\text{right}}$ were sampled from $\mathcal{N}(\textbf{0},\textbf{I})$, and in HaP $y_{\text{left}}$ and $y_{\text{right}}$ are each the normal label of a real encoded image (from Algorithm 1), chosen so that they have the same orientation. For CelebA we used the $d=100$ model and for HaP we used $d=50$ (chosen after consulting the IS and FID results across all four $d$ values in Appendix Table 3 while preferring a higher dimension ($d\in\{50,100\}$) to retain more visual detail).